# Distribution functions for the modelling of accretion remnants in Milky Way-like galaxies: insights from IllustrisTNG

### v1

@@ Abstract
pubyear: 2024pagerange: Distribution functions for the modelling of accretion remnants in Milky Way-like galaxies: insights from IllustrisTNGDistribution functions for the modelling of accretion remnants in Milky Way-like galaxies: insights from IllustrisTNG
@@ 1 Introduction
In this work we turn our attention to distribution functions (DFs), perhaps the least-used technique of the aforementioned for modelling individual remnants. Density profiles have been fit to the whole stellar halo for some time (for contemporary examples see Deason et al., 2019; Mackereth & Bovy, 2020), and have recently been fit specifically to GS/E (Han et al., 2022; Lane et al., 2023). Multi-component Gaussian models, representing simple DFs, have been applied to kinematic data in order to study GS/E as well with good success (e.g. Lancaster et al., 2019; Fattahi et al., 2019; Iorio & Belokurov, 2021). More recently, Lane et al. (2022) and Lane et al. (2023) employed simple constant-anisotropy DFs to study and fit density profiles to GS/E, however they did not actively fit the DF to data but constructed a DF based on assumptions about GS/E to aid in sample selection. The active fitting of DFs to data represents a next step in the study of major accretion remnants such as GS/E, but also smaller remnants and in-situ structures.
@@
In this paper we will identify and study accretion remnants in the IllustrisTNG cosmological simulation (Nelson et al., 2019a), and investigate how well commonly used classes of DFs are able to fit the data. Our aim is for this work to provide a basis from which future studies of remnants using DFs can proceed confidently. The paper is laid out as follows: § 2 provides an overview of the IllustrisTNG simulations and describes the selection of Milky Way analogs as well and major mergers. § 3 describes the DFs that we use and provides the fitting methodology. In § 4 we assess how well each DF model is matched to the data. We end with a discussion and conclusion in sections § 5 and § 6 respectively.
@@ 3.1 The constant-anisotropy and Osipkov-Merritt distribution functions
where $\sigma_{[\theta,\phi,r]}$ are the spherical velocity dispersions in the polar, azimuthal, and radial directions respectively. $\beta$ has the domain $(-\infty,1]$, with negative values corresponding to a tangentially biased system, a value of 0 corresponding to an isotropic or ergodic (i.e. the DF depends only on energy) system, and positive values corresponding to a radially biased system. Finally, the function $f_{1}$ relates the tracer density of the DF to the underlying potential. For ergodic ($\beta=0$) DFs $f_{1}$ can be solved by a simple integral inversion (Eddington, 1916), however for non-zero $\beta$ the calculation is more involved. For more information see Cuddeford (1991), An & Evans (2006), chapter 4.3 of Binney & Tremaine (2008), and Appendix A of Lane et al. (2022) which specifically describes the implementation of this DF we use in galpy (Bovy, 2015).
@@ 3.3 Fitting distribution functions to data
The procedure to fit the constant anisotropy and both standard and superposition Osipkov-Merritt DFs is similar. We consider that the required ingredients to construct these DFs are the gravitational potential $\Phi$, the density profile of the tracer $\rho$, and a parameter which specifies the anisotropy: $\beta$ for the constant-anisotropy DF, $r_{a}$ for the Osipkov-Merritt DF, and $\{r_{a,1},r_{a,2},k_{\mathrm{om}}\}$ for the superposition Osipkov-Merritt DF. As mentioned above, these models are computationally expensive to work with we therefore build the DF from each of these individual components, which are tractable to obtain on their own, as opposed to fitting the entire model to data all at once.
@@ 4.1 Two case studies: GS/E and the Sequoia
To select a GS/E analog we examine remnants with $\beta>0.8$ (numbering 6 total), and choose one with stellar mass of $\approx 6\times 10^{9}$ $\mathrm{M}_{\odot}$ and anisotropy of $\beta=0.88$, which was deposited in a 6:1 stellar mass ratio merger (30:1 dark matter mass ratio) at $z\approx 0.8$. In choosing the GS/E analog we focus most on anisotropy, attempting to match the value for GS/E of $\sim 0.9$. The stellar mass of GS/E has been subject to a wide range of estimates, spanning nearly two decades from $10^{8}$ to $10^{10}$ $\mathrm{M}_{\odot}$, however more direct measurements of the density profile of GS/E have tended to settle on a lower stellar mass range of $1.5-7.2\times 10^{8}$ $\mathrm{M}_{\odot}$ (Mackereth & Bovy, 2020; Han et al., 2022; Lane et al., 2023). These estimates typically correspond with a total mass ratio for the merger in the range of 1:4-1:10 (normally assuming a GS/E merger epoch of $z\approx 2$). So the mass ratios are in broad agreement with estimates for GS/E, while the total stellar mass is on the high-end of estimates. The accretion epoch for GS/E is thought to be 7-11 Gyr ago ($z\sim 0.8-2.4$), and so this merger is on the very lowest end of that range. This partially explains the higher total stellar mass. Only one of the candidate remnants with high-anisotropy merged at $z>1$, however it only has a present-day stellar mass of only $\sim 10^{7}$ $\mathrm{M}_{\odot}$, much too low to be considered an appropriate analog for GS/E.
@@ 5.2.1 The constant anisotropy and Osipkov-Merritt DFs
But frustratingly none of the DF models were able to fully reproduce the observed morphology of either the GS/E or Sequoia analog remnants in the energy-angular momentum space. While each model was able to do a reasonable job of getting the locus and extent in energy and angular momentum of the N-body particles, they were unable to capture some of the semantic details in the distributions, such as the specific trends in the width and location of the distribution of angular momentum with varying energy. These failures highlight that while these models can be effective they are clearly limited. Coupling this with the fact that the DFs — particularly the superposition Osipkov-Merritt DF – can match the dispersions and anisotropy of the remnants well suggests that higher order moments of the DF could play an important role.
With this in mind it does appear to be the case that high-anisotropy remnants are better represented using the Osipkov-Merritt (and therefore naturally the superposition Osipkov-Merritt as well) model than the constant-anisotropy DF. The constant-anisotropy model substantially overpredicts the velocity anisotropy via the radial velocity dispersion at small radii. This leads to an overprediction of the density of stars with low angular momenta, as is seen in Figure 6. These observations are grounded in the mathematical behaviour of the constant-anisotropy models, which are known to be unstable for orbits with low tangential velocity (see Binney, 2014).
@@ 6 Summary and conclusions
    We find clear evidence that highly anisotropy stellar populations with $\beta>0.8$, such as GS/E, are better modelled using the standard Osipkov-Merritt DF as opposed to the constant-anisotropy DF; an even better model is the superposition of two Osipkov-Merritt DFs. While these remnants may have high average anisotropy, the profiles nearly always become more ergodic in the center of the Galaxy. We estimate that an Osipkov-Merritt profile with a scale radius of 2-4 kpc would provide a reasonable DF model for GS/E.
@@ Acknowledgements
JMML and JB acknowledge financial support from NSERC (funding reference number RGPIN-2020-04712). We are very greatful to Ted Mackereth, who helped to conceptualize this project.

### v2

@@ Abstract
pubyear: 2025pagerange: Distribution functions for the modelling of accretion remnants in Milky Way-like galaxies: insights from IllustrisTNGDistribution functions for the modelling of accretion remnants in Milky Way-like galaxies: insights from IllustrisTNG
@@ 1 Introduction
In this work we turn our attention to distribution functions (DFs), perhaps the least-used technique of the aforementioned for modelling individual remnants. Density profiles have been fit to the whole stellar halo for some time (for contemporary examples see Deason et al., 2019; Mackereth & Bovy, 2020), and have recently been fit specifically to GS/E (Han et al., 2022; Lane et al., 2023). Multi-component Gaussian models, representing simple DFs, have been applied to kinematic data in order to study GS/E as well with good success (e.g. Lancaster et al., 2019; Fattahi et al., 2019; Iorio & Belokurov, 2021). More recently, Lane et al. (2022) and Lane et al. (2023) employed simple constant-anisotropy DFs to study and fit density profiles to GS/E, however they did not actively fit the DF to data but constructed a DF based on assumptions about GS/E to aid in sample selection. It has become clear in the Gaia era that remnants in the stellar halo have unique spatial, kinematic, and chemical properties, making DFs better suited than just density profiles owing to their ability to fold in kinematic information implicitly as well as chemistry with simple additions (e.g. Sanders & Binney, 2015b). Furthermore, there are many DFs, including those that we propose to study in this work, which are better physically motivated than simpler models such as Gaussians, which are often preferred due to their flexibility and ubiquity. DFs have now been used to study many other dynamical groups in the Milky Way, including the stellar disk, bar, bulge, as well as the stellar halo as a whole, in addition to the dark halo. We believe that they can now also be applied to better study, classify, and understand the individual merger remnants that comprise the stellar halo as well. In particular, active fitting of DFs to data represents a next step in the study of major accretion remnants such as GS/E, but also smaller remnants and in-situ structures.
@@
In this paper we will identify and study accretion remnants in the IllustrisTNG cosmological simulation (Nelson et al., 2019a), and investigate how well commonly used classes of DFs are able to fit the data. We focus specifically on a small number of well known DFs in this work aware that there are other forms and novel techniques that may be studied in the future, and will provide some discussion of these avenues at the end of the paper. Our overall aim for this work is to extract insight using these simple DFs to inform future work on the topic, but more importantly to provide a basis from which upcoming studies of remnants using DFs can start to proceed more confidently. We do not aim to provide definitive techniques for optimal DF fits to simulated data, but hope that this study may lay the ground work to accomplish this goal down the line. The paper is laid out as follows: § 2 provides an overview of the IllustrisTNG simulations and describes the selection of Milky Way analogs as well and major mergers. § 3 describes the DFs that we use and provides the fitting methodology. In § 4 we assess how well each DF model is matched to the data. We end with a discussion and conclusion in sections § 5 and § 6 respectively.
@@ 3.1 The constant-anisotropy and Osipkov-Merritt distribution functions
where $\sigma_{[\theta,\phi,r]}$ are the spherical velocity dispersions in the polar, azimuthal, and radial directions respectively. $\beta$ has the domain $(-\infty,1]$, with negative values corresponding to a tangentially biased system, a value of 0 corresponding to an isotropic or ergodic (i.e. the DF depends only on energy) system, and positive values corresponding to a radially biased system. Finally, the function $f_{1}$ relates the tracer density of the DF to the underlying potential. For ergodic ($\beta=0$) DFs $f_{1}$ can be solved by a simple integral inversion (Eddington, 1916), however for non-zero $\beta$ the calculation is more involved. For more information see Cuddeford (1991), An & Evans (2006a), Evans & An (2006), chapter 4.3 of Binney & Tremaine (2008), and Appendix A of Lane et al. (2022) which specifically describes the implementation of this DF we use in galpy (Bovy, 2015).
@@ 3.3 Fitting distribution functions to data
The procedure to fit the constant anisotropy and both standard and superposition Osipkov-Merritt DFs is similar. We consider that the required ingredients to construct these DFs are the gravitational potential $\Phi$, the density profile of the tracer $\rho$, and a parameter which specifies the anisotropy: $\beta$ for the constant-anisotropy DF, $r_{a}$ for the Osipkov-Merritt DF, and $\{r_{a,1},r_{a,2},k_{\mathrm{om}}\}$ for the superposition Osipkov-Merritt DF. The choice to fit DFs in this way, using the density profile, underlying potential, and anisotropy-specifying parameters, is not the only way to accomplish this task. Since we have energy and angular momenta for the remnant particles we could fit the DF directly in the space of $(\mathcal{E},L)$ using either a custom form or the constant anisotropy and Osipkov-Merritt DFs. While this is possible, by focusing on the density profile and anisotropy profile—as we do—we are grounding the computed DFs in those observables, which we argue are just as important if not more important than the specific distribution in ($\mathcal{E},L$). DFs computed in this way are guaranteed to reproduce (at least to be very close, see below) the target density profile and anisotropy parameters. Additionally, while we do not attempt to prove this, we suspect that to fit the DF piece-wise as we do would produce very similar results to a DF fitted to integrals of motion. Finally it is relevent that if we were to fit the DF in $(\mathcal{E},L)$ space using the constant anisotropy or Osipkov-Merritt forms then it would add significant computational expense to the fitting process, as computation of such DFs is time consuming.
Note that, as derived by An & Evans (2006b), there is a limit on the value of $\beta$ in the inner part of an anisotropic DF such that $\beta$ must be less than half the inner power law slope of the target density profile. This limit is applicable for self-consistent DFs, and is modified by a small factor when the potential and density and independent as is the case for the DFs used here. We recognize this limitation and navigate it in practice by setting any negative regions of computed DFs to 0 (see Lane et al., 2022, for more information on how this is done during construction of the DF). Such a choice means that any constructed DF will have density and anisotropy profile slightly different than are specified by the input ingredients, but the compromise ensures the physicality of the DF. Additionally, we only ever work with DFs at radii larger than the softening length (0.29 kpc) which slightly mitigates the impact of this effect.
Finally, here we neglect triaxiality in both the remnants and the potential. The host potentials are certainly expected to be flattened, and potentially triaxial in a cosmological context. In the inner parts of the system the baryonic disk flattens the potential, and the dark matter halo, which dominates the outskirts, is also expected to be flattened and mildly triaxial (Dubinski, 1994; Abadi et al., 2010). In order to assess the triaxiality of our sample we compute the moment of inertia tensor for each remmant. Diagonalizing the tensor gives the eigenvalues, which are the squares of the axis scale lengths $(a>b>c)$. We assess the distributions of $b/a$ and $c/a$, finding them peaked near their respective medians of 0.91 and 0.84. The distribution of $b/a$ ranges between 0.8 and 1, while $c/a$ ranges between 0.7 and 0.95.
So we see that most remnants are triaxial to a small or modest degree, indicating that a spherical assumption in our analysis is reasonable. Additionally, proper treatment of triaxiality is challenging. This is primarily because of the lack of good, general DFs with independent density profiles and potentials, but also because obtaining reliable integrals of motion for such systems is challenging. So here we work with stricly spherical models, and leave triaxiality to follow up studies. Additional discussion of the impact of this choice will be presented in § 5.2.2.
@@ 4.1 Two case studies: GS/E and the Sequoia
To select a GS/E analog we examine remnants with $\beta>0.8$ (numbering 6 total), and choose one with stellar mass of $\approx 6\times 10^{9}$ $\mathrm{M}_{\odot}$ and anisotropy of $\beta=0.88$, which was deposited in a 6:1 stellar mass ratio merger (30:1 dark matter mass ratio) at $z\approx 0.8$. In choosing the GS/E analog we focus most on anisotropy, attempting to match the value for GS/E of $\sim 0.9$. The stellar mass of GS/E has been subject to a wide range of estimates, spanning nearly two decades from $10^{8}$ to $10^{10}$ $\mathrm{M}_{\odot}$, however more direct measurements of the density profile of GS/E have tended to settle on a lower stellar mass range of $1.5-7.2\times 10^{8}$ $\mathrm{M}_{\odot}$ (Mackereth & Bovy, 2020; Han et al., 2022; Lane et al., 2023). These estimates typically correspond with a total mass ratio for the merger in the range of 1:4–1:10 (normally assuming a GS/E merger epoch of $z\approx 2$). So the mass ratios are in broad agreement with estimates for GS/E, while the total stellar mass is on the high-end of estimates. The accretion epoch for GS/E is thought to be 7-11 Gyr ago ($z\sim 0.8-2.4$), and so this merger is on the very lowest end of that range. This partially explains the higher total stellar mass. Only one of the candidate remnants with high-anisotropy merged at $z>1$, however it only has a present-day stellar mass of only $\sim 10^{7}$ $\mathrm{M}_{\odot}$, much too low to be considered an appropriate analog for GS/E.
@@ 5.2.1 The constant anisotropy and Osipkov-Merritt DFs
But frustratingly none of the DF models were able to fully reproduce the observed morphology of either the GS/E or Sequoia analog remnants in the energy-angular momentum space. While each model was able to do a reasonable job of getting the locus and extent in energy and angular momentum of the N-body particles, they were unable to capture some of the semantic details in the distributions, such as the specific trends in the width and location of the distribution of angular momentum with varying energy. These failures highlight that while these models can be effective they are clearly limited. Coupling this with the fact that the DFs — particularly the superposition Osipkov-Merritt DF — can match the dispersions and anisotropy of the remnants well suggests that higher order moments of the DF could play an important role.
With this in mind it does appear to be the case that high-anisotropy remnants are better represented using the Osipkov-Merritt (and therefore naturally the superposition Osipkov-Merritt as well) model than the constant-anisotropy DF. The constant-anisotropy model substantially overpredicts the velocity anisotropy via the radial velocity dispersion at small radii. This leads to an overprediction of the density of stars with low angular momenta, as is seen in Figure 6.
@@ 5.2.2 Consideration for other DFs
In this study we focus on fitting the density profile, underlying potential, and anisotropy profile of each remnant to create our DFs. A different approach could be to focus exclusively on energy and angular momentum (in addition possible to other integrals of motion) and attempt to find functional forms that describe their distributions. These would be valid DFs, being functions of integrals of motion, but would not likely correspond to density profiles and potentials of the functional forms we are familiar with (e.g. the two-power models we use in this study). But these DFs would be useful in that they would, by design, be well matched to integrals of motion distributions, which are the common spaces used to study the Milky Way stellar halo at present. This type of strategy may be best suited as a tailored approach to studying specific remnants such as GS/E in the Milky Way stellar halo, since it may not be possible to find functional forms that can be broadly applied. Yet, it is not clear that the DFs we use in this study are of broad use either, and so working solely in integrals of motion space may ultimately prove just as effective if not better.
Finally, in this work we neglect triaxiality both in the underlying potential and the density profile for each remnant. The potential of the Milky Way is certainly flattened in the inner parts of the Galaxy thanks to the stellar disk, and in the outer parts of the Galaxy the potential-governing dark matter halo may be oblate and mildly triaxial (Law & Majewski, 2010; Vera-Ciro & Helmi, 2013; Bovy et al., 2016). The physical geometry of most remnants in the stellar halo has yet to be thoroughly studied and it is likely that most are at least somewhat flattened, following the general shape of the Milky Way potential, and some may be triaxial. The only remnant to have had its 3D shape thoroughly assessed is GS/E, which is found to be modestly triaxial (Iorio & Belokurov, 2021; Han et al., 2022; Lane et al., 2023). Certainly in the broader cosmological context it is expected that halos with a baryonic component are triaxial (Dubinski, 1994; Abadi et al., 2010). Is it therefore important to consider triaxiality among remnants and the host potential in future studies that seek to use DFs.
The action-based DFs noted above can incorporate flattening by treating the total angular momentum $L$ differently from the vertical component $L_{\mathrm{z}}$. Additionally, Sanders & Evans (2015) demonstrated the construction of action-based, triaxial density profiles with self-consistent DFs based on those of Williams & Evans (2015). But the problem of the general applicability of triaxial DFs is threefold. The first is the models add a great deal of complexity to the problem at hand, both in computational load and the possible necessity for fine tuning. Second, in so far as the models are based on actions they are not always grounded in physical observables, and the ability of the models to match key observables may be limited. Third, the models can be limited by the accessibility of integrals of motion (or their lack of existence). Integrals of motion for a triaxial system can be acquired using approximate (Sanders & Binney, 2015a) or perturbative techniques (Binney, 2018), or in a more tailored, accurate manner for systems such as the Galactic bar (Binney, 2020), but such approaches are not applicable broadly. The impact of our choice to work only with spherical systems is challenging to gauge, but unlikely to be substantial. Certainly the DFs will not as faithfully reproduce the underlying data as if they did include a triaxial perscription, however, as demonstrated above many observables are well described, in a broad sense by our spherical DFs. It could be the case that some of the discrepencies in the energy and angular momenta examined in § 4.1 could be caused by the triaxiality of the underlying potential, the density profile of the remnant, or both. Nonetheless we believe that future work should be done to study the applicability of triaxial DFs for the modelling of remnants.
@@ 5.2.3 The DF for GS/E
Building on the discussion from the previous section, while we are confident that Osipkov-Merritt DFs are more suitable for anisotropic remnants like GS/E compared with constant anisotropy models, we do not claim to have generated a definitive DF for the remnant. It is certainly important to consider that GS/E has been demonstrated to exhibit triaxiality (Han et al., 2022; Lane et al., 2023), indicating that a more accurate DF may be one of the action-based forms including a triaxial perscription. Furthermore, it is an open question whether or not GS/E is sufficiently phase-mixed to be considered in equilibrium. On one hand the majority of the remnant occupies the inner Milky Way stellar halo ($<20~{}\mathrm{kpc}$) where the dynamical times are $\lesssim 1~{}\mathrm{Gyr}$, much shorter than the supposed $8-10$ Gyr age of the remnant (Montalbán et al., 2021). On the other hand the initial evidence of the triaxiality of the remnant from Simion et al. (2019) and Iorio & Belokurov (2019)—that it appears linked with the Hercules-Aquila cloud and Virgo Overdensity—may also hint that it is not in a relaxed state if these are non-equilibrium structures instead of simply being concentrations along one triaxial axis. Finally, the evidence presented by Chandra et al. (2023) of GS/E-affiliated substructure in the outter stellar halo ($>60$ kpc) clearly indicates that the debris is not in equilibrium at those radii. Indeed, a good candidate DF for GS/E, in combination with stellar survey data with a well-understood selection function, could possibly be used to assess whether GS/E is in equilibrium in the inner Milky Way. Regardless, further work is certainly needed to create a definitive DF for GS/E which includes triaxiality and can accurately match other observables such as the anisotropy, velocity dispersions, and density profile across a large range of radii.
@@ 6 Summary and conclusions
    We find clear evidence that highly anisotropy stellar populations with $\beta>0.8$, such as GS/E, are better modelled using the standard Osipkov-Merritt DF as opposed to the constant-anisotropy DF; an even better model is the superposition of two Osipkov-Merritt DFs. While these remnants may have high average anisotropy, the profiles nearly always become more ergodic in the center of the Galaxy. We estimate that an Osipkov-Merritt profile with a scale radius of 2–4 kpc would provide a reasonable DF model for GS/E.
@@ Acknowledgements
We thank the referee for their comments, which have certainly improved the quality of the manuscript. JMML and JB acknowledge financial support from NSERC (funding reference number RGPIN-2020-04712). We are very greatful to Ted Mackereth, who helped to conceptualize this project.